# Non-minimally Coupled Running Curvaton: A Unified Approach to Early-Universe Inflation and Phantom Dark Energy

### v2

@@ Abstract
[…]
@@ I Introduction
This observational hint poses a significant theoretical challenge. Canonical single-field quintessence models are typically restricted to the region $w\geq-1$ Caldwell et al. (1998); Zlatev et al. (1999), while naive attempts to realize phantom behavior often lead to pathologies such as ghost degrees of freedom or gradient instabilities Vikman (2005); Hu (2005); Caldwell et al. (2003). Consequently, constructing a consistent theoretical framework that explains both the early inflationary epoch and the specific late-time phantom dynamics favored by DESI—without introducing instabilities—remains an open problem Creminelli et al. (2009); Feng et al. (2005); Nojiri et al. (2005). However, this issue can be addressed by introducing a non-minimal coupling Yao et al. (2025); Wolf et al. (2026); Tiwari et al. (2025); Sánchez López et al. (2025); Wang et al. (2025); Tita et al. (2025); Wolf et al. (2025); Adam et al. (2025). Although numerous approaches exist to study evolving dark energy—including the Quintom scenario Cai et al. (2025); Yang et al. (2025, 2024), $f(R)$ gravitational models Nojiri et al. (2025b, a); Odintsov and Oikonomou (2026), even using the minimally coupling to investigate the dynamical DE Gialamas et al. (2025b) that represents the local effects of DE Gialamas et al. (2025a). We restrict our attention in this paper to the non-minimally coupled dark energy model.
@@
The “Running Curvaton” model Liu and Xu (2020a) was previously proposed as a unified approach to connect early-universe physics with late-time acceleration. In this framework, the curvaton field generates primordial perturbations during inflation and survives to serve as the dark energy candidate today. However, in its original minimally coupled form, the model behaves as a thawing quintessence field, which is theoretically unable to access the phantom regime (w<1w<-1) suggested by the new data.
In this paper, we propose an extension of the Running Curvaton model by introducing a non-minimal gravitational coupling of the form $\xi\chi^{2}R$. We demonstrate that this modification provides the necessary geometric corrections to the effective energy-momentum tensor in the Jordan frame, enabling the equation of state to cross the phantom divide naturally Faraoni (2000); Elizalde et al. (2004); Gannouji et al. (2006). This extension allows the model to populate the $w_{0}-w_{a}$ observational plane favored by DESI 2024, resolving the tension faced by the standard model.
@@
A critical concern when modifying a unified model is whether the new physics spoils successful predictions for the early universe. We address this by deriving a parameter re-tuning scheme. We show that the geometric correction to the effective mass of the curvaton during inflation can be exactly compensated by shifting the coupling parameters. As a result, the standard predictions for the spectral index ($n_{s}$) and non-Gaussianity ($f_{NL}\approx 5/4r_{dec}$) are strictly preserved, ensuring consistency with Planck observations Akrami and others (2020).
@@
The paper is organized as follows: Section II outlines the non-minimally coupled Running Curvaton model. Section III demonstrates the preservation of early-universe perturbation predictions via parameter re-tuning. Section IV analyzes the late-time evolution, deriving the effective equation of state and confronting the model with DESI constraints and stability requirements. Finally, Section V concludes with a discussion of our findings.
@@ II The Model
The second part of the curvaton potential (2) is distinct from the original version, $V(\chi)=\frac{1}{2}\frac{g}{M_{P}^{2}}V(\phi)\chi^{2}+\lambda_{0}e^{-\lambda_{1}\chi/M_{P}}$ (where $\lambda_{0}\sim V_{1}$). We modify the potential to the form in Eq. (2) because it provides a better fit to DESI observations Gu and others (2025); Abdul Karim and others (2025), which allows for a dynamic equation of state that better captures the $w_{0}-w_{a}$ evolution suggested by the DESI 2024 data. While the exponential potential in the original model reproduces thawing quintessence (TQ) Wolf and Ferreira (2023), its CPL parameters ($w_{0},w_{a}$) are in tension with, or only marginally consistent with, the most recent constraints Wolf et al. (2024).
@@ IV.1 Stability Analysis: Sound Speed
This result ($c_{s}^{2}=1>0$) confirms that despite the effective equation of state crossing into the phantom regime ($w_{\rm eff}<-1$) due to the non-minimal coupling $\xi\chi^{2}R$, the scalar field perturbations propagate at the speed of light and are free from gradient instabilities. The phantom behavior in our model is a frame-dependent effect arising from the modified Einstein equations, rather than a pathological kinetic term.
In the following subsection, we will constrain the parameter via the Parameterized Post-Newtonian (PPN) approximation.
IV.2 Screening Mechanism at Solar System Scales
@@ IV.3 Modified Friedmann Equations
In a flat FLRW universe, the variation of Action (1) yields the modified Friedmann equations (Detailed calcultion can be found in Appendix D). Neglecting the inflaton field at late times, we have:
@@ IV.5 Crossing the Phantom Divide
Figure 1: The evolution for the equation of state $w(z)$. The blue solid line corresponds $(\xi,\lambda,V_{0},V_{1})=(2.3,0.75,0.94,5.29)$. The green dashed-dotted line is $(\xi,\lambda,V_{0},V_{1})=(2.1,1.42,0.72,2.24)$. The red dotted line is $(\xi,\lambda,V_{0},V_{1})=(3.0,0.5,0.9,5.0)$. The orange dashed line is $(\xi,\lambda,V_{0},V_{1})=(1.5,0.82,0.65,2.67)$. We have set $M_{P}=1$ and $\rho_{M,0}=3M_{P}^{2}\Omega_{m}^{0}=0.93$ with $\Omega_{m}^{0}=0.31$ and $H_{0}=69\rm~km~s^{-1}~Mpc^{-1}$.
@@
IV.6 Constraints from DESI
In this section, we will show our model will be consistent with DESI Gu and others (2025). In our case, $w_{0}$ denotes the present value of $w_{\rm eff}|_{a=1}$. Based on (27) and (28), we could derive
@@ IV.6 Constraints from DESI
Figure 2: Observational Constraints vs. Non-minimal Coupling Curvaton. This figure compares the theoretical predictions of our non-minimally coupled Running Curvaton model with the 68% and 95% confidence levels from DESI Y5, Pantheon+, and Union3 datasets. The four distinct models correspond to the four cases in Fig. 1. The black star corresponds to the $\Lambda CDM$.
@@
Figure 3: $68\%$ and $95\%$ C.L. posterior distributions for runing curvaton model from from the combination of DESI BAO ++ Pantheon, obtained from sampling uniformly in the parameters via fig. 2.
@@
To constrain the free parameters of the non-minimal coupling model, namely $\theta=\{\xi,\lambda,V_{0},V_{1}\}$, we performed a Monte Carlo parameter scan using the rejection sampling technique. We assumed uniform priors for all parameters within physically motivated ranges: $\xi\in[0,5]$, $\lambda\in[0,2]$, $V_{0}\in[0.5,1.5]$, and $V_{1}\in[1.0,6.0]$. For each randomly sampled set of parameters, we numerically solved the background equations of motion (Eq. 32 and Eq. 33) to derive the corresponding dark energy equation-of-state parameters at present day, $w_{0}$, and its evolution, $w_{a}$. These theoretical predictions were then mapped onto the observational $w_{0}-w_{a}$ plane. We utilized the $95\%$ confidence level (C.L.) contours obtained from the combination of DESI DR2 BAO, Pantheon+ supernovae, and Planck CMB data as our acceptance criterion. Specifically, we adopted a top-hat likelihood approach: a parameter set was accepted only if its predicted $(w_{0},w_{a})$ fell within the observational $95\%$ C.L. region; otherwise, it was rejected. Fig. 3 shows the resulting marginalized posterior distributions and 2D confidence regions for the model parameters, derived from $20,000$ Monte Carlo trials. Based on the results shown in this figure, the best-fit parameters are $(\xi,\lambda,V_{0},V_{1})=(2.9,0.96,0.76,3.5)$. We observe that Model B lies almost entirely outside the DESI constraints, as its $\xi$ value is smaller than the minimum threshold of $\xi=1.6$ shown in Fig. 2. In contrast, Model C aligns well with the best-fit region in Fig. 3. Overall, the results in Fig. 3 demonstrate strong consistency with those presented in Fig. 2.
@@ V Discussion and Conclusion
The most significant achievement of this model is its ability to naturally explain the dynamical dark energy evolution favored by DESI 2024 combined with CMB and SNIa data. We showed that the interplay between the non-minimal coupling $\xi\chi^{2}R$ and the plateau-like potential $V(\chi)$ generates necessary geometric corrections to the effective energy-momentum tensor. This mechanism drives the effective equation of state $w_{eff}$ to cross the phantom divide (w<1w<-1) in the past and evolve towards the quintessence regime (w>1w>-1) today, precisely populating the observational contour of $w_{0}>-1$ and $w_{a}<0$. In light of the most recent DESI 2025 data Gu and others (2025), Fig. 2 demonstrates that our model falls well within the observational confidence contours. Additionally, the corresponding best-fit model parameters are illustrated in Fig. 3.
@@
In summary, the Non-minimally Coupled Running Curvaton offers a theoretically sound and observationally consistent pathway to unify the physics of the early and late universe. It not only alleviates the current tensions in the dark energy sector but also provides a concrete mechanism for dynamical phantom crossing without violating fundamental stability conditions. Future high-precision surveys, such as Euclid and LSST, will further constrain the derivative of the equation of state, offering a decisive test for the specific $w_{0}-w_{a}$ trajectory predicted by this framework. Furthermore, this work could be extended into multi-field frameworks Liu and Prokopec (2021); Zhang et al. (2023) and $f(R)$ gravity theories Liu (2018); Liu et al. (2018, 2021). The running curvaton model also provides a mechanism for generating primordial black holes (PBHs) as candidates for dark matter Liu (2023). Additionally, one could investigate the effects of non-minimal coupling on PBH formation.

### v3

# Non-minimally Coupled Running Curvaton for DESI-preferred Dynamical Dark Energy and Hubble Tension

@@ Abstract
Recent DESI 2025 results have renewed interest in dynamical dark energy models with phantom crossing. In this work, we extend the running curvaton scenario by introducing a non-minimal gravitational coupling, $\xi\chi^{2}R$. We show that this coupling can drive the late-time equation of state into the DESI-preferred region in the $(w_{0},w_{a})$ plane. We further find that, at leading order, the early-universe predictions of the original running curvaton model can be preserved through a parameter redefinition, while the scalar sector remains free of gradient instability in the regime considered. A parameter scan identifies viable regions and a benchmark solution that also shifts the background-inferred value of $H_{0}$ upward. These results suggest that the non-minimally coupled running curvaton provides a viable framework for DESI-motivated dark energy evolution, with possible implications for the Hubble tension.
@@ I Introduction
Specifically, the SH0ES collaboration reported a local value of $H_{0}=73.04\pm 1.04$ km/s/Mpc Riess et al. (2022), leading to a tension exceeding $5\sigma$. This discrepancy is further reinforced by independent late-time probes, including strong gravitational lensing from the H0LiCOW/TDCOSMO collaboration Wong et al. (2020) and the TRGB (Tip of the Red Giant Branch) method Freedman (2021), although some debates on systematic uncertainties remain. This ”Early vs. Late” dichotomy strongly suggests the need for physics beyond the standard paradigm, potentially involving either early-time modifications to the sound horizon Poulin et al. (2019) or late-time dynamical dark energy evolution. Most notably, recent results from the Dark Energy Spectroscopic Instrument (DESI) 2025, when combined with Supernovae Type Ia (SNIa) and CMB datasets, favor a dynamical dark energy component over a cosmological constant Gu et al. (2025); Wang (2024). Specifically, the data prefers an equation of state that crosses the so-called “phantom divide” (w<1w<-1) at late times, characterized by the parameter space $w_{0}>-1$ and $w_{a}<0$ Gu et al. (2025); Rubin et al. (2023); Brout et al. (2022); Abbott et al. (2024); Colgáin et al. (2024); Pan and Ye (2025).
@@
This observational hint poses a significant theoretical challenge. Canonical single-field quintessence models are typically restricted to the region $w\geq-1$ Caldwell et al. (1998); Zlatev et al. (1999), while naive attempts to realize phantom behavior often lead to pathologies such as ghost degrees of freedom or gradient instabilities Vikman (2005); Hu (2005); Caldwell et al. (2003). Consequently, constructing a consistent theoretical framework that explains both the early inflationary epoch and the specific late-time phantom dynamics favored by DESI—without introducing instabilities—remains an open problem Creminelli et al. (2009); Feng et al. (2005); Nojiri et al. (2005). However, this issue can be addressed by introducing a non-minimal coupling Yao et al. (2025); Wolf et al. (2026); Tiwari et al. (2025); Sánchez López et al. (2025); Wang et al. (2025); Tita et al. (2025); Wolf et al. (2025); Adam et al. (2025). Although numerous approaches exist to study evolving dark energy—including the Quintom scenario Cai et al. (2025); Yang et al. (2025, 2024), $f(R)$ gravitational models Nojiri et al. (2025a, b); Odintsov and Oikonomou (2026), even using the minimally coupling to investigate the dynamical DE Gialamas et al. (2025a) that represents the local effects of DE Gialamas et al. (2025b). We restrict our attention in this paper to the non-minimally coupled dark energy model.
Under the framework of non-minimally coupling, there are some models for investigating the Hubble tension. The $f(R)$ gravity could account for this problem in the Jordan frame Schiavone et al. (2023); Montani et al. (2026, 2023); Fazzari et al. (2025). Additionally, Ref. Barroso Varela and Bertolami (2024) demonstrated that the non-minimally coupling between the curvature and gravity could confront with the CMB, without the traditional cosmological constant. Meanwhile, this puzzle could also be released within $2\sigma$ in the Horndeski gravity Tiwari et al. (2024). The most popular method for explaining the evolving dark energy is the Quintessence Sánchez López et al. (2025), that could also address this problem using the newest DESI data.
@@
The “Running Curvaton” model Liu and Xu (2020a) was previously proposed as a unified approach to connect early-universe physics with late-time acceleration. In this framework, the curvaton field generates primordial perturbations during inflation and survives to serve as the dark energy candidate today. However, in its original minimally coupled form, the model behaves as a thawing quintessence field, which is theoretically unable to access the phantom regime (w<1w<-1) suggested by the new data. In this paper, we propose an extension of the Running Curvaton model by introducing a non-minimal gravitational coupling of the form $\xi\chi^{2}R$. We demonstrate that this modification provides the necessary geometric corrections to the effective energy-momentum tensor in the Jordan frame, enabling the equation of state to cross the phantom divide naturally Faraoni (2000); Elizalde et al. (2004); Gannouji et al. (2006). This extension allows the model to populate the $w_{0}-w_{a}$ observational plane favored by DESI 2025, resolving the tension faced by the standard model.
@@
A critical concern when modifying a unified model is whether the new physics spoils successful predictions for the early universe. We address this by deriving a parameter re-tuning scheme. We show that the geometric correction to the effective mass of the curvaton during inflation can be exactly compensated by shifting the coupling parameters. As a result, the standard predictions for the spectral index ($n_{s}$) and non-Gaussianity ($f_{NL}\approx 5/4r_{dec}$) are strictly preserved, ensuring consistency with Planck observations Akrami et al. (2020).
@@
The paper is organized as follows: Section II outlines the non-minimally coupled Running Curvaton model. Section III demonstrates the preservation of early-universe perturbation predictions via parameter re-tuning. Section IV analyzes the late-time evolution, deriving the effective equation of state and confronting the model with DESI constraints and stability requirements, meanwhile it will automatically account for the Hubble tension. Finally, Section V concludes with a discussion of our findings.
@@ II The Model
Motivation for the modified potential.
The choice of the late-time potential in Eq. (2) is motivated by phenomenology rather than by an attempt to claim a full statistical fit. In the original minimally coupled Running Curvaton setup, the pure exponential potential leads to late-time dynamics similar to thawing quintessenc Wolf and Ferreira (2023), whose trajectories in the $(w_{0},w_{a})$ plane are in tension with, or at best only marginally consistent with, the region preferred by recent DESI data Gu et al. (2025); Abdul Karim et al. (2025). By contrast, the modified form
$V(\chi)=V_{1}\left(1-V_{0}e^{-\lambda\chi/M_{P}}\right)$
introduces additional shape freedom that allows the field evolution to interpolate more efficiently between a nearly frozen stage and a later rolling phase. When combined with the non-minimal coupling $\xi\chi^{2}R$, this structure more readily generates trajectories exhibiting the past phantom crossing behavior favored by the DESI-motivated region with $w_{0}>-1$ and $w_{a}<0$ Wolf et al. (2024). Therefore, the potential in Eq. (2) should be understood as a phenomenologically motivated extension that facilitates access to the desired late-time dynamics in our parameter scan, rather than as a uniquely selected or fully fitted form.
@@ IV.1 Stability Analysis: Sound Speed
This result ($c_{s}^{2}=1>0$) confirms that despite the effective equation of state crossing into the phantom regime ($w_{\rm eff}<-1$) due to the non-minimal coupling $\xi\chi^{2}R$, the scalar field perturbations propagate at the speed of light and are free from gradient instabilities. The phantom behavior in our model is a frame-dependent effect arising from the modified Einstein equations, rather than a pathological kinetic term. In addition to the positivity of $c_{s}^{2}$, viability of the model also requires the effective Planck mass squared, $M_{\rm eff}=M_{P}^{2}-\xi R^{2}$, to remain positive throughout the cosmological evolution considered here. Futhermore, the tensor propagation speed remains luminal $c_{T}^{2}=1$ due to $G_{5}=0$ and $G_{4,X}=0$, in agreement with current gravitational-wave constraints.
IV.2 Remarks on Local Screening and Possible Solar-System Viability
@@
Here, we emphasize that this argument suggests a possible suppression of fifth-force effects in high-density environments. A quantitative confrontation with Solar-System tests, however, would require a dedicated analysis of the static spherically symmetric solution and the corresponding post-Newtonian parameters, which lies beyond the scope of the present work.
@@ IV.3 Modified Friedmann Equations
In a flat FLRW universe, the variation of Action (1) yields the modified Friedmann equations (Detailed calculation can be found in Appendix D). Neglecting the inflaton field at late times, we have:
@@ IV.5 Crossing the Phantom Divide
Figure 1: Evolution of the effective dark-energy equation of state $w(z)$ for four representative parameter choices in the non-minimally coupled running curvaton model. The curves illustrate that increasing $\xi$ facilitates access to a phantom phase at intermediate redshift, while the trajectories approach w=1w=-1 as $\xi$ approaches zero.
@@
IV.6 Comparison with DESI-preferred Regions
Our goal in this section is not to perform a full global fit, but rather to identify regions of parameter space in which the model can reproduce the DESI-preferred late-time dark-energy behavior.
In this section, we will show our model will be consistent with DESI Gu et al. (2025). In our case, $w_{0}$ denotes the present value of $w_{\rm eff}|_{a=1}$. Based on (27) and (28), we could derive
@@ IV.6 Comparison with DESI-preferred Regions
Figure 2: Representative model trajectories in the $w_{0}-w_{a}$ plane compared with the DESI-motivated confidence contours shown for reference. The colored markers correspond to the four benchmark parameter choices used in Fig. 1. This figure is intended as a visual comparison.
@@
Figure 3: Distributions of accepted parameter samples for the non-minimally coupled running curvaton model, obtained from the parameter scan described in Sec. IV F. The darker and lighter shaded regions indicate the regions enclosing $68\%$ and $95\%$ of the accepted samples, respectively.
@@
To constrain the free parameters of the non-minimal coupling model, namely $\theta=\{\xi,\lambda,V_{0},V_{1}\}$, we performed a Monte Carlo parameter scan using the rejection sampling technique. We assumed uniform priors for all parameters within physically motivated ranges: $\xi\in[0,5]$, $\lambda\in[0,2]$, $V_{0}\in[0.5,1.5]$, and $V_{1}\in[1.0,6.0]$. For each randomly sampled set of parameters, we numerically solved the background equations of motion (Eq. 32 and Eq. 33) to derive the corresponding dark energy equation-of-state parameters at present day, $w_{0}$, and its evolution, $w_{a}$. These theoretical predictions were then mapped onto the observational $w_{0}-w_{a}$ plane. We utilized the $95\%$ confidence level (C.L.) contours obtained from the combination of DESI DR2 BAO, Pantheon+ supernovae, and Planck CMB data as our acceptance criterion. Specifically, we adopted a top-hat likelihood approach: a parameter set was accepted only if its predicted $(w_{0},w_{a})$ fell within the observational $95\%$ C.L. region; otherwise, it was rejected. Fig. 3 shows the resulting distributions of accepted samples and 2D confidence regions for the model parameters, derived from $20,000$ Monte Carlo trials. Based on the results shown in this figure, the best-fit parameters are $(\xi,\lambda,V_{0},V_{1})=(2.9,0.96,0.76,3.5)$. We observe that Model B lies almost entirely outside the DESI constraints, as its $\xi$ value is smaller than the minimum threshold of $\xi=1.6$ shown in Fig. 2. In contrast, Model C aligns well with the best-fit region in Fig. 3. Overall, the results in Fig. 3 demonstrate strong consistency with those presented in Fig. 2.
IV.7 Implications for the Hubble Tension
Besides fitting the DESI-preferred dynamical dark energy region in the $w_{0}$$w_{a}$ plane, the late-time phantom crossing in the non-minimally coupled Running Curvaton model also has a direct geometric implication for the Hubble tension. The key point is that the non-minimal coupling modifies the late-time expansion history while leaving the pre-recombination physics essentially unchanged through the re-tuning mechanism
$g_{0}^{\rm obs}=g_{0}+2\xi.$ (34)
As discussed in Sec. III, this guarantees that the primordial perturbation sector is preserved, so that the standard early-universe quantities, including the comoving sound horizon at recombination, remain unchanged to a very good approximation.
IV.7.1 CMB acoustic-scale consistency
The strongest early-universe constraint comes from the precisely measured CMB acoustic scale,
$\theta_{*}=\frac{r_{s}(z_{*})}{D_{M}(z_{*})},$ (35)
where $r_{s}(z_{*})$ is the comoving sound horizon at recombination and $D_{M}(z_{*})$ is the comoving angular-diameter distance to the last-scattering surface. Because the re-tuning mechanism preserves the pre-recombination expansion history in our setup, $r_{s}(z_{*})$ is fixed at essentially the same value as in the standard scenario. Therefore, consistency with the observed $\theta_{*}$ requires that
$D_{M}(z_{*})=c\int_{0}^{z_{*}}\frac{dz}{H(z)}$ (36)
remains unchanged.
It is convenient to rewrite the Hubble parameter in the form
$H(z)=H_{0}E(z),$ (37)
with
$E^{2}(z)=\Omega_{m,0}(1+z)^{3}+\Omega_{{\rm DE},0}^{\rm eff}\,f(z),$ (38)
where the dark-energy evolution factor is
$f(z)=\exp\left[3\int_{0}^{z}\frac{1+w_{\rm eff}(z^{\prime})}{1+z^{\prime}}\,dz^{\prime}\right].$ (39)
Hence,
$D_{M}(z_{*})=\frac{c}{H_{0}}\int_{0}^{z_{*}}\frac{dz}{E(z)}.$ (40)
IV.7.2 Late-time phantom crossing and the shift of $H_{0}$
[…]
$w_{\rm eff}(z)<-1.$ (41)
In this regime,
$1+w_{\rm eff}(z)<0,$ (42)
and therefore the exponent in $f(z)$ becomes negative, implying
$f_{\rm phantom}(z)<1$ (43)
for the relevant redshift interval. As a consequence, the normalized expansion history satisfies
$E_{\rm phantom}(z)<E_{\Lambda{\rm CDM}}(z)$ (44)
at intermediate redshifts where the phantom behavior is active.
[…]
$\int_{0}^{z_{*}}\frac{dz}{E(z)}$ (45)
increases. To keep the CMB distance $D_{M}(z_{*})$ fixed and preserve the observed acoustic scale $\theta_{*}$, the prefactor $H_{0}^{-1}$ must decrease. Equivalently, the inferred present-day Hubble constant must increase:
$H_{0}^{\rm phantom}>H_{0}^{\Lambda{\rm CDM}}.$ (46)
Therefore, the same late-time phantom-crossing dynamics that improves the fit to the DESI data naturally shifts the preferred value of $H_{0}$ upward, partially alleviating the tension between the CMB-inferred and local distance-ladder determinations.
IV.7.3 Numerical implication for the Hubble constant
To quantify the geometric effect discussed above, we numerically integrate the background evolution from $z=0$ to $z\approx z_{*}\simeq 1090$, while imposing the condition that the comoving angular diameter distance to the last-scattering surface remains fixed. Since the re-tuning mechanism preserves the early-universe physics and keeps the sound horizon $r_{s}(z_{*})$ unchanged, the shift in the inferred Hubble constant is entirely induced by the modified late-time expansion history.
For the benchmark solution near the DESI best-fit region, namely
$(\xi,\lambda,V_{0},V_{1})=(2.9,\,0.96,\,0.76,\,3.5),$ (47)
which corresponds to Model C in Fig. 2 and Fig. 3, the numerical integration shows that the preferred Hubble constant is shifted upward to
$H_{0}\approx 73.94~{\rm km\,s^{-1}\,Mpc^{-1}}.$ (48)
[…]
Fig. 4 illustrates this result explicitly. The black dashed curve and red solid curve represent the Gaussian distributions corresponding to the Planck 2018 and SH0ES 2022 determinations, respectively, while the green dash-dotted curve denotes the theoretical prediction of our non-minimally coupled Running Curvaton model for the benchmark parameter set. One can clearly see that the geometric phantom-crossing effect shifts the model prediction from the Planck value toward the SH0ES region, thus providing a viable late-time mechanism for reconciling the two measurements.
[…]
IV.7.4 Remarks on interpretation
We stress that, at the present stage, the resolution of the Hubble tension discussed here should be understood as a background-level geometric mechanism: the late-time phantom evolution lowers $E(z)$ over part of cosmic history, which in turn pushes the CMB-compatible value of $H_{0}$ upward. A complete statistical assessment requires a joint likelihood analysis using CMB, BAO, and supernova data, together with the exact background solution for the non-minimally coupled curvaton field. Nevertheless, Nevertheless, the qualitative mechanism is clear at the background level and follows directly from the modified late-time expansion history induced by the coupling $\xi\chi^{2}R$.
@@ V Discussion and Conclusion
The most significant achievement of this model is its ability to naturally explain the dynamical dark energy evolution favored by DESI 2025 combined with CMB and SNIa data. We showed that the interplay between the non-minimal coupling $\xi\chi^{2}R$ and the plateau-like potential $V(\chi)$ generates necessary geometric corrections to the effective energy-momentum tensor. This mechanism drives the effective equation of state $w_{eff}$ to cross the phantom divide (w<1w<-1) in the past and evolve towards the quintessence regime (w>1w>-1) today, precisely populating the observational contour of $w_{0}>-1$ and $w_{a}<0$. In light of the most recent DESI 2025 data Gu et al. (2025), Fig. 2 demonstrates that our model falls well within the observational confidence contours. Additionally, the corresponding best-fit model parameters are illustrated in Fig. 3.
(b). Implication to the Hubble tension
The late-time phantom crossing driven by $\xi\chi^{2}R$ modifies the expansion history by lowering the normalized expansion rate $E(z)$ as shown in Sec. IV.7. To maintain the precisely measured CMB acoustic scale $\theta_{*}$, the inferred Hubble constant is naturally shifted upward to $H_{0}\approx 73.94\text{ km s}^{-1}\text{Mpc}^{-1}$. This result falls within the $1\sigma$ region of the SH0ES 2022 local distance-ladder measurement, indicating that the model has the potential to ease the tension at the level of background geometry.
@@
In summary, the Non-minimally Coupled Running Curvaton offers a theoretically sound and observationally consistent pathway to unify the physics of the early and late universe. It not only alleviates the current tensions in the dark energy sector but also provides a concrete mechanism for dynamical phantom crossing and Hubble tension without violating fundamental stability conditions. Future high-precision surveys, such as Euclid and LSST, will further constrain the derivative of the equation of state, offering a decisive test for the specific $w_{0}-w_{a}$ trajectory predicted by this framework. Furthermore, this work could be extended into multi-field frameworks Liu and Prokopec (2021); Zhang et al. (2023) and $f(R)$ gravity theories Liu (2018); Liu et al. (2018, 2021). The running curvaton model also provides a mechanism for generating primordial black holes (PBHs) as candidates for dark matter Liu (2023). Additionally, one could investigate the effects of non-minimal coupling on PBH formation.